OpenStudy (anonymous):

Without dividing, do you expect five-eighths divided by one-sixteenth to be greater than or less than 1? Explain.

OpenStudy (anonymous):

@fabsam14 Five eighths = 5/8 One-sixteenth = 1/16 When you divide any number A by any number B, then the result is less than 1 if A is lesser than B, and greater than 1 when A is greater than B. We know that 5/8 is close to half, since 5 is close to 4 and we can write the fraction as 4/8, which equals 1/2. We know that 1/16 is close to 0, since 1 is close to 0 and we can write the fraction as 0/16. Therefore 5/8 is greater than 1/16. Thus we can say that the result of dividing 5/8 by 1/16 will be greater than 1.